OpenStudy (anonymous):

The denominator of the fraction 4/15 is increased by an amount so that the value of the resulting fraction is 1/5. Determine the amount by which the denominator was increased.

OpenStudy (anonymous):

4/15 -> 3/15 -1/15 is the increase. Where did you go from there?

OpenStudy (anonymous):

4/15 + a = 1/5 multiply both sides by 15: 4 + 15a = 1 15a = -3 a = -3/15 a = -1/5

OpenStudy (lalaly):

4+15a=3 15a=-1 a=-1/15